A tennis ball is launched straight upward with an initial velocity of 24.5 m/s from the edge of a cliff that is 117.6 meters above the ground. Which quadratic equation could be used to correctly determine when the ball will hit the ground:
4.9t2 + 24.5t + 117.6 = 0
-4.9t2 - 24.5t + 117.6 = 0
-4.9t2 + 24.5t - 117.6 = 0
4.9t2 + 24.5t - 117.6 = 0
-4.9t2 + 24.5t + 117.6 = 0
Thank you

Answer by Edwin McCravy(20055)

The formula from physics is

h(t) = h0 + v0t - 4.9t2

v0 = initial velocity = 12.5 m/s
h0 = initial height = 117.6 m
h(t) = final height = 0 m

Substituting:

h(t) = h0 + v0t - 4.9t2

   0 = 117.6 + 24.5t - 4.9t2

Can you tell which one is equivalent to that?

Edwin